OpenStudy (anonymous):

I have one more question @campbell_st and I was wondering if I do the same thing we did last time.

OpenStudy (campbell_st):

if you post it i'll have a look

OpenStudy (campbell_st):

Ok, this is a teapezium or trapezoid, depending on where you live the formula is \[A = \frac{1}{2} \times (a + b) \times h\] a and b are the lengths of the parallel sides and h is the perpendicular or right angled, distance between the 2 lines so what do you think a = ? b = ? and h = ?

OpenStudy (anonymous):

I think a and b are 8 and 11 and h is 4.

OpenStudy (anonymous):

@campbell_st I got 38. Am I correct?

OpenStudy (campbell_st):

thats correct.